OpenStudy (anonymous):

Solve the differential equation dy/dt = y^2 / (4-t^2)^(1/2) 1/sin-1(t) + c 1/sin-1(t/2) + c 2/sin-1(t/2) + c 2/tan-1(t/2) + c

OpenStudy (kc_kennylau):

Put the terms with y on one side and terms with t on another.

OpenStudy (anonymous):

Ok, then what? dy/y^2=dt/(4-t^2)^(1/2)

OpenStudy (kc_kennylau):

integrate both sides

OpenStudy (kc_kennylau):

\[\int\frac{dy}{y^2}=\int\frac{dt}{\sqrt{4-t^2}}\]

OpenStudy (anonymous):

that's not an answer though?

OpenStudy (kc_kennylau):

I'd never give you the answer

OpenStudy (kc_kennylau):

You have to find it out yourself

OpenStudy (anonymous):

I'm confused. If you integrate both sides, then you'd get sin-1(t/2). Which isn't a choice?

OpenStudy (kc_kennylau):

Because the left hand side is still not calculated

OpenStudy (anonymous):

oh! okay. i see. the left side would be -1/y. so it'd be -1/y=arcsin(t/2) + c you multiple both sides by y then divide both sides by arcsin(t/2) + c. so you get -1/arcsin(t/2) + c ...but that's still not answer choice, right?

OpenStudy (kc_kennylau):

well,,,

OpenStudy (anonymous):

did i do something wrong?

OpenStudy (kc_kennylau):

Don't think so